Hydrated strontium chloride has a molar mass of 266.6 A student heats 5.332g of SrCl2
1
Expert's answer
2020-12-08T05:34:41-0500

M (SrCl2•X H2O) = 266.6 g/mol

Moles whole thing = 5.332 / 266.6 = 0.02 moles

H2O lost: (5.332 - 3.892) / 18 = 0.08 moles

0.08 / 0.02 = 4 so 4 moles of water were lost

2 moles of water left.

So the product is SrCl2 • 2H2O